ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 49, 35 Migórski, Stanisław, Jagiellonian University of Cracow, Poland Optimal control of hemivariational inequalities Optimal control and optimal shape design problems for systems governed by elliptic and parabolic hemivariational inequalities (being a generalization of variational inequalities) are considered. The goal is to establish the existence of optimal solutions for problems with integral cost functionals. In the shape design problems the cost depend on the shape through the solution of the hemivariational inequality defined on the domain which should be optimized. The results are directly applicable to many mechanical problems involving nonmonotone, nonconvex and possible multivalued relations between fluxes and forces. Reference: Z. Denkowski, and S. Migorski, Optimal shape design problems for a class of systems described by hemivariational inequalities, J. Global Optimiz. 12 (1998), 37–59.